A 6-month promissory note dated April 12, 2020 is made at 5.2% for $2,800. What is the present value of the note 40 days later if money is worth 7%? (Note: Assume there are 183 days of interest on 6 months here.)
1
Expert's answer
2020-04-20T15:01:31-0400

1.Determine the future value of the bill:

"FV=PV(1+\frac{183}{366}\times i)"


"FV=2800(1+\frac{183}{366}\times0.052)"


"FV=2800\times1.026=2872.8"

2."P=2872.8(1-\frac{183-40}{366}\times0.07)=2872.8\times0.973=2794.23"